​Southeastern Bell stocks a certain switch connector at its central warehouse for supplying field service offices.

Southeastern Bell stocks a certain switch connector at its central warehouse for supplying field service offices. The yearly demand for these connectors is 15,000 units Southeastern estimates its annual holding cost for this item to be $24 per unit. The cost to place and process an order from the supplier is $77. The company operates 300 days per year, and the lead time to receive an order from the supplier is 3 working days.


a)Find the economic order quantity.

b) Find the annual holding costs.

c) Find the annual ordering costs.

d) What is the reorder point?
Answer #1

a) EOQ = sqroot{(2*Annual Demand*Ordering cost)/Holding cost}

= sqroot{(2*15000*77)/24}

= 310 Approximately

b) Annual Holding cost = Average inventory * Holding cost per unit

= (310/2)*24

= $3720

c) Annual Ordering cost = Number of orders * Ordering cost

= (15000/310)*77

= $3726

d) Reorder Point = Average Daily Demand * Lead Time

= (15000/300)*3

= 150 Units